OpenStudy (vera_ewing):

Point B lies on AC between A and C. AB=3x-8 and BC=5x+3 . Find AC.

OpenStudy (vera_ewing):

@amistre64 @thomaster @whpalmer4 @aaronq help please! i have to learn this by today!

OpenStudy (amistre64):

the sum of the parts is equal to the whole ....

OpenStudy (amistre64):

AC = AB + BC of course

OpenStudy (vera_ewing):

uh oh how do i do that?

OpenStudy (amistre64):

addition

OpenStudy (vera_ewing):

so 3x-8 + 5x+3 ???

OpenStudy (amistre64):

correct

OpenStudy (vera_ewing):

ok I got 8x-3

OpenStudy (amistre64):

close, what is -8+3?

OpenStudy (vera_ewing):

-5

OpenStudy (amistre64):

then lets go with 8x - 5

OpenStudy (vera_ewing):

so that is the answer?

OpenStudy (amistre64):

that is what AB is equal to, without knowing anything else, then yes. that is a simplified result.

OpenStudy (vera_ewing):

so AC=8x-5 ?

OpenStudy (amistre64):

AC = AB + BC AB = BC = 8x - 5 therefore AC = 8x - 5, if we have to solve for x, then we work some algebra to get x all alone

OpenStudy (amistre64):

AC = AB + BC AB + BC = 8x - 5 ** edited the typo therefore AC = 8x - 5, if we have to solve for x, then we work some algebra to get x all alone

OpenStudy (vera_ewing):

ok thanks! LH bisects GLI . What is the measure of KLH ?

OpenStudy (amistre64):

i have no idea what objects you are referring to ....

OpenStudy (vera_ewing):

there is a picture....

OpenStudy (amistre64):

pictures help :)

OpenStudy (vera_ewing):



OpenStudy (vera_ewing):

the bottom 3 points are G, L, and K in that order

OpenStudy (amistre64):

yeah, youd have to type and reclick a few times, it doesnt like linear text for some reason

OpenStudy (amistre64):